Balliol College Oxford University Oxford OX1 3BJ

September 19, 2000

Sticky Compsoc Network Somewhere in Trinity

Dear Sticky,

Are you enjoying your semi-retirement? For no reason here’s an equation:

\frac{sinx}{x} = 2

and some more random text.

Best Wishes, Tom Ford